OpenStudy (anonymous):

How large of a sample size is required to estimate the mean within 1.6 units with 90% confidence, given s=0.47.

OpenStudy (anonymous):

The answer is exactly 1, but I don't understand how

OpenStudy (anonymous):

didn't learn this sry cant help

ganeshie8 (ganeshie8):

use below : Margin of error= \(\large Z^* \frac{s}{\sqrt{n}} \)

ganeshie8 (ganeshie8):

you're given : Margin of error = \(1.6\) \(s = 0.47\)

ganeshie8 (ganeshie8):

you need to find \(Z^*\) corresponding to 90% CI and plugin above

OpenStudy (anonymous):

yes I get 0.611229, but its asking how large of a sample, my zscore table says 1.645

ganeshie8 (ganeshie8):

http://www.math.upenn.edu/~chhays/zscoretable.pdf

ganeshie8 (ganeshie8):

yes, for 90% CI, \(Z^* = 1.65\)

ganeshie8 (ganeshie8):

plugin that value and solve \(n\)

ganeshie8 (ganeshie8):

\(\large 1.6 = 1.65 \frac{0.47}{\sqrt{n}}\)

ganeshie8 (ganeshie8):

solve \(n\)

OpenStudy (anonymous):

0.23350036035

ganeshie8 (ganeshie8):

yes, so the sample size needs to be a MINIMUM of 0.2335..

ganeshie8 (ganeshie8):

think a bit, can u really sample 0.2335... persons ?

OpenStudy (anonymous):

ohhh bless your heart, the problem is simpified!

OpenStudy (anonymous):

i get it, thank you

ganeshie8 (ganeshie8):

np... u wlc :)